# Efficient Classical Algorithms for Simulating Gaussian Boson Sampling on Graphs

### v1

@@ C.1 Complete Bipartite Graph $G_{m,n}$
For convenience, let $T^{+}(k):=\{(M,M+e)\mid\text{$M$ is a matching of size $k$}\}$ and $T^{-}(k):=\{(M,M-e)\mid\text{$M$ is a matching of size $k$}\}$. The critical observation is that for complete bipartite graphs, our construction ensures that in one set $T^{+}(k)$ (or $T^{-}(k)$), each transition shares identical congestion values. Notice that the original construction in [16] only considers ordering the vertices in $V$, and obtains a unique ordering of components by the smallest vertex. The start vertex is defined as the smallest vertex in a cycle, and the smaller of the two endpoints in a path. This construction is not sufficient to ensure the identical congestion requirement.
@@
$\min\{|I|,|F|\}-1\leq|\gamma_{I,F}^{(i)}|\leq\max\{|I|,|F|\}+1.$
Figure 5 shows an example of one canonical path from $I$ to $F$.
Figure 5: A possible canonical path between matchings $I$ and $F$.
@@
$\Pr[M\to M\cup\{e\}]$ $=\frac{1}{mn}\cdot\frac{\lambda^{2}}{1+\lambda^{2}},$
$\Pr[M\cup\{e\}\to M]$ $=\frac{1}{mn}\cdot\frac{\operatorname{Haf}(G_{M})}{\operatorname{Haf}(G_{M\cup\{e\}})}\cdot\frac{1}{1+\lambda^{2}}\geq\frac{1}{mn}\cdot\frac{1}{|M|+1}\cdot\frac{1}{1+\lambda^{2}}.$
@@
$\varrho_{t}:=\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{(I,F,i)\in\mathrm{cp}(t)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|.$
@@
$N_{p}=\binom{n}{p}\binom{m}{p}\cdot p!,\quad N_{p,q}=\begin{cases}N_{p}\cdot(n-p)(m-p)&\text{ if }q=p+1;\\ N_{p}\cdot p&\text{ if }q=p-1.\end{cases}$
We first consider $t\in T^{+}(k)$. Suppose $t=(M,M^{\prime})$ with $|M|=k$ and $|M^{\prime}|=k+1$, note that by Proposition 17, if the size of $I$ and the size of $F$ are either both less than $k-1$ or both greater than $k+1$, then any path from $I$ to $F$ cannot pass through any transition in the set $T^{+}(k)$. Therefore, we have
$\quad N_{k,k+1}\cdot\varrho_{t}=$ $\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{t^{\prime}=(M^{\prime},M^{\prime}+e^{\prime}):|M|=|M^{\prime}|}\sum_{(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|$
$\leq$ $\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{t^{\prime}=(M^{\prime},M^{\prime}+e^{\prime}):|M|=|M^{\prime}|}\bigg{(}\sum_{|I|\geq k-1,|F|\leq k+1}\sum_{i:(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|$
$+\sum_{|I|\leq k+1,|F|\geq k-1}\sum_{i:(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|\bigg{)}$
$=$ $\frac{1}{\pi(M)P(M,M^{\prime})}\bigg{(}\sum_{|I|\geq k-1,|F|\leq k+1}|\gamma_{IF}|\sum_{i=1}^{l(I,F)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|$
$+\sum_{|I|\leq k+1,|F|\geq k-1}|\gamma_{IF}|\sum_{i=1}^{l(I,F)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}|\bigg{)}$
$\leq$ $\frac{1}{\pi(M)P(M,M^{\prime})}\bigg{(}\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}|^{2}+\sum_{I:|I|\geq k-1}\sum_{F:|F|\leq k+1}\pi(I)\pi(F)|\gamma_{IF}|^{2}\bigg{)}$
$=$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}|^{2}$
$\leq$ $2\cdot\frac{mn(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot N_{k,k+1}\cdot\frac{1}{N_{k,k+1}}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$
Notice that $\frac{N_{l+1}\cdot\lambda^{2(l+1)}(l+1)!}{N_{l}\cdot\lambda^{2l}l!}=\lambda^{2}(n-l)(m-l)\geq 4\lambda^{2}$ for $l<n-1$, when $\lambda>\frac{1}{2}$, we have
$\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k,k+1}\lambda^{2k}k!}=$ $\frac{1}{(n-k)(m-k)}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k+1}\cdot\lambda^{2(k+1)}(k+1)!}\cdot\frac{N_{k+1}\cdot\lambda^{2}(k+1)}{N_{k}}$
$\leq$ $\frac{1}{(n-k)(m-k)}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\cdot\lambda^{2}(n-k)(m-k)$
$=$ $\lambda^{2}\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right).$
@@
$N_{k,k-1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}|^{2}$
$\leq$ $2\cdot mn^{2}(1+\lambda^{2})\cdot(2n)^{2}\cdot N_{k,k-1}\cdot\frac{1}{N_{k,k-1}}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$
@@
$\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k,k-1}\lambda^{2k}k!}=\frac{1}{k}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k}\cdot\lambda^{2k}k!}\leq\frac{1}{k}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\leq\frac{4\lambda^{2}}{4\lambda^{2}-1}+1.$
@@
$\varrho_{t}$ $\leq\max\left\{2\cdot\frac{mn(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot\lambda^{2}\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right),2\cdot mn^{2}(1+\lambda^{2})\cdot(2n)^{2}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\right\}$
$=8mn^{4}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}.$
Let $\bar{\lambda}=\max\{1/\lambda,\lambda\}$, noting that $\lambda^{2|M|}\cdot|M|!\leq\bar{\lambda}^{2n}\cdot n!$ for any matching $M$, and the number of matchings in the complete bipartite graph $G$ is bounded by $\frac{m!}{(m-n)!}\cdot n!\leq(m!)^{2}\leq m^{2m}$. We have the crude bound $\ln\pi(M)^{-1}\leq\ln{\left({\bar{\lambda}}^{2n}\cdot n!\cdot m^{2m}\right)}\leq 2n\ln|{\lambda}|+2m\ln{m}$, which holds uniformly over all matchings of $G$, by Theorem 10, we establish a polynomial mixing time of double-loop Glauber on complete bipartite graphs:
Lemma 18.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}/r_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(16mn^{4}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)\right)^{-1}$
in Lemma 4. Then the failure probability of inner layer sampling in each iteration can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma 18, after $r_{\frac{\epsilon}{2}}$ iterations, the total variance can be bounded by $\frac{\epsilon}{2}+\eta_{\frac{\epsilon}{2}}\cdot r_{\frac{\epsilon}{2}}=\epsilon$. Therefore, we give a polynomial-time upper bound for the mixing of double-loop Glauber on complete bipartite graph $G_{m,n}$:
Theorem 19.
@@
C.2 Dense Bipartite Graphs
In this section, we extend the results from Appendix C.1 to non-complete bipartite graphs, providing proofs of Theorem 2 and Theorem 6. A natural approach is to leverage the aforementioned congestion calculations. We discuss dense bipartite graphs $G$: there exists a constant $\xi$, such that each vertex is non-adjacent to at most $\xi$ vertices in the opposite partition.
@@ C.2 Dense Bipartite Graphs
$\pi(M)P(M,M^{\prime})=\frac{1}{|E|}\cdot\frac{\lambda^{2}}{1+\lambda^{2}}\cdot\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M}),$
we can rewrite (15) as follows:
ϱ:=assignitalic-$\varrho:=$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\pi(I)\pi(F)|\gamma_{IF}|\bigg{\}}$
$=$ $\max_{t=(M,M^{\prime})}\bigg{\{}|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\cdot\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\bigg{\}}.$ (16)
For any bipartite graph $G$, let $G^{\prime}$ be the corresponding complete bipartite graph with the same vertex set as $G$. Let $\varrho_{t}^{\prime}$ denote the congestion of transition $t=(M,M^{\prime})$ in $G^{\prime}$, $E^{\prime}$ denote the edge set of $G^{\prime}$, $P^{\prime}(M,M^{\prime})$ denote the corresponding transition probability of double-loop Glauber for $G^{\prime}$, and $\operatorname{Haf}^{\prime}(M)$ denote the corresponding hafnain of the subgraph induced by matching $M$ in $G^{\prime}$, which is equal to $|M|!$. We have already established an estimate for $\varrho_{t}^{\prime}$ in Appendix C.1, and now we can use this estimate to bound $\varrho_{t}$ in $G$ by estimate $\operatorname{Haf}(M)$ via $\operatorname{Haf}^{\prime}(M)$ for any matching $M$.
@@
$1\geq\frac{\operatorname{Haf}(M)}{\operatorname{Haf}^{\prime}(M)}\geq\frac{(n^{\prime}-\xi)!}{n^{\prime}!}\geq\frac{1}{n^{\prime\xi}}\geq\frac{1}{n^{\xi}}.$
@@
ϱ=italic-$\varrho=$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\cdot|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg{\}}$
$\leq$ $\max_{t=(M,M^{\prime})}\bigg{\{}\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}^{\prime}(I)\lambda^{2|F|}\operatorname{Haf}^{\prime}(F)|\gamma_{IF}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}^{\prime}(\tilde{M})n^{-\xi}\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}^{\prime}(X)n^{-\xi}}\cdot|E^{\prime}|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg{\}}$
$\leq$ $n^{2\xi}\varrho^{\prime}.$
@@
By the analysis of Lemma 18, we have the congestion $\varrho^{\prime}$ of double-loop Glauber on complete bipartite graph $G_{m,n}$ is bounded by $\varrho^{\prime}=8mn^{4}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}$. Then we have the congestion $\varrho$ of double-loop Glauber on bipartite graph $G$ is bounded by $\varrho\leq n^{2\xi}\cdot\varrho^{\prime}=8mn^{4+2\xi}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}$. Thus again by Theorem 10, the mixing time of double-loop Glauber on complete bipartite graph $G_{m,n}$ is less than $16mn^{4+2\xi}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)$. ∎
Furthermore, we complete the proof of Theorem 6 by extending the analysis of Theorem 19 to the case of dense bipartite graphs.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(16mn^{4+2\xi}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)\right)^{-1}$
in Lemma 4. Then the failure probability of each inner layer sampling step can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma 18, after
$16mn^{4+2\xi}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)$

### v2

@@ A Complete Bipartite Graph $G_{m,n}$
For convenience, let $T^{+}(k):=\{(M,M+e)\mid\text{$M$ is a matching of size $k$}\}$ and $T^{-}(k):=\{(M,M-e)\mid$
$\text{$M$ is a matching of size $k$}\}$. The critical observation is that for complete bipartite graphs, our construction ensures that in one set $T^{+}(k)$ (or $T^{-}(k)$), each transition shares identical congestion values. Notice that the original construction in [18] only considers ordering the vertices in $V$, and obtains a unique ordering of components by the smallest vertex. The start vertex is defined as the smallest vertex in a cycle, and the smaller of the two endpoints in a path. This construction is not sufficient to ensure the identical congestion requirement.
@@
$\min\{|I|,|F|\}-1\leq|\gamma_{I,F}^{(i)}|\leq\max\{|I|,|F|\}+1.$ (S31)
Supplementary Fig. 1 shows an example of one canonical path from $I$ to $F$.
Supplementary Fig. 1: A possible canonical path between matchings $I$ and $F$.
@@
$\Pr[M\to M\cup\{e\}]$ $=\frac{1}{mn}\cdot\frac{\lambda^{2}}{1+\lambda^{2}},$ (S32)
$\Pr[M\cup\{e\}\to M]$ $=\frac{1}{mn}\cdot\frac{\operatorname{Haf}(G_{M})}{\operatorname{Haf}(G_{M\cup\{e\}})}\cdot\frac{1}{1+\lambda^{2}}\geq\frac{1}{mn}\cdot\frac{1}{|M|+1}\cdot\frac{1}{1+\lambda^{2}}.$ (S33)
@@
$\varrho_{t}:=\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{(I,F,i)\in\mathrm{cp}(t)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|.$ (S34)
@@
[…]
We first consider $t\in T^{+}(k)$. Suppose $t=(M,M^{\prime})$ with $|M|=k$ and $|M^{\prime}|=k+1$, note that by Proposition S2, if the size of $I$ and the size of $F$ are either both less than k1k-1 or both greater than $k+1$, then any path from $I$ to $F$ cannot pass through any transition in the set $T^{+}(k)$. Therefore, we have
$\quad N_{k,k+1}\cdot\varrho_{t}=$ $\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{t^{\prime}=(M^{\prime},M^{\prime}+e^{\prime}):|M|=|M^{\prime}|}\sum_{(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|$ (S36)
$\leq$ $\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{t^{\prime}=(M^{\prime},M^{\prime}+e^{\prime}):|M|=|M^{\prime}|}\bigg(\sum_{|I|\geq k-1,|F|\leq k+1}\sum_{i:(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|$
$+\sum_{|I|\leq k+1,|F|\geq k-1}\sum_{i:(I,F,i)\in\mathrm{cp}(t^{\prime})}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|\bigg)$ (S37)
$=$ $\frac{1}{\pi(M)P(M,M^{\prime})}\bigg(\sum_{|I|\geq k-1,|F|\leq k+1}|\gamma_{IF}^{(i)}|\sum_{i=1}^{l(I,F)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|$
$+\sum_{|I|\leq k+1,|F|\geq k-1}|\gamma_{IF}^{(i)}|\sum_{i=1}^{l(I,F)}p^{(i)}(I,F)\pi(I)\pi(F)|\gamma_{IF}^{(i)}|\bigg)$ (S38)
$\leq$ $\frac{1}{\pi(M)P(M,M^{\prime})}\bigg(\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}+\sum_{I:|I|\geq k-1}\sum_{F:|F|\leq k+1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}\bigg)$ (S39)
$=$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k+1}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}$ (S40)
$\leq$ $2\cdot\frac{mn(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot N_{k,k+1}\cdot\frac{1}{N_{k,k+1}}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$ (S41)
Notice that $\frac{N_{l+1}\cdot\lambda^{2(l+1)}(l+1)!}{N_{l}\cdot\lambda^{2l}l!}=\lambda^{2}(n-l)(m-l)\geq 4\lambda^{2}$ for l<n1l<n-1, when $\lambda>\frac{1}{2}$, we have
$\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k,k+1}\lambda^{2k}k!}=$ $\frac{1}{(n-k)(m-k)}\cdot\frac{\sum_{l=1}^{k+1}N_{l}\cdot\lambda^{2l}l!}{N_{k+1}\cdot\lambda^{2(k+1)}(k+1)!}\cdot\frac{N_{k+1}\cdot\lambda^{2}(k+1)}{N_{k}}$ (S42)
$\leq$ $\frac{1}{(n-k)(m-k)}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\cdot\lambda^{2}(n-k)(m-k)$ (S43)
$=$ $\lambda^{2}\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right).$ (S44)
@@
$N_{k,k-1}\cdot\varrho_{t}\leq$ $2\cdot\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{I:|I|\leq k}\sum_{F:|F|\geq k-1}\pi(I)\pi(F)|\gamma_{IF}^{(i)}|^{2}$ (S45)
$\leq$ $2\cdot mn^{2}(1+\lambda^{2})\cdot(2n)^{2}\cdot N_{k,k-1}\cdot\frac{1}{N_{k,k-1}}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!\cdot\sum_{l=k-1}^{n}N_{l}\cdot\lambda^{2l}l!}{\lambda^{2k}k!\cdot\sum_{l=1}^{n}N_{l}\cdot\lambda^{2l}l!}.$ (S46)
@@
$\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k,k-1}\lambda^{2k}k!}=\frac{1}{k}\cdot\frac{\sum_{l=1}^{k}N_{l}\cdot\lambda^{2l}l!}{N_{k}\cdot\lambda^{2k}k!}\leq\frac{1}{k}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\leq\frac{4\lambda^{2}}{4\lambda^{2}-1}+1.$ (S47)
@@
$\varrho_{t}$ $\leq\max\left\{2\cdot\frac{mn(1+\lambda^{2})}{\lambda^{2}}\cdot(2n)^{2}\cdot\lambda^{2}\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right),2\cdot mn^{2}(1+\lambda^{2})\cdot(2n)^{2}\cdot\left(\frac{4\lambda^{2}}{4\lambda^{2}-1}+1\right)\right\}$ (S48)
$=8mn^{4}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}.$ (S49)
Let $\bar{\lambda}=\max\{1/\lambda,\lambda\}$, noting that $\lambda^{2|M|}\cdot|M|!\leq\bar{\lambda}^{2n}\cdot n!$ for any matching $M$, and the number of matchings in the complete bipartite graph $G$ is bounded by $\frac{m!}{(m-n)!}\cdot n!\leq(m!)^{2}\leq m^{2m}$. We have the crude bound $\ln\pi(M)^{-1}\leq\ln{\left({\bar{\lambda}}^{2n}\cdot n!\cdot m^{2m}\right)}\leq 2n\ln|{\lambda}|+2m\ln{m}$, which holds uniformly over all matchings of $G$, by Theorem S2, we establish a polynomial mixing time of double-loop Glauber on complete bipartite graphs:
Lemma S6.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}/r_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(16mn^{4}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)\right)^{-1}$ (S50)
in Lemma 1. Then the failure probability of inner layer sampling in each iteration can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma S6, after $r_{\frac{\epsilon}{2}}$ iterations, the total variance can be bounded by $\frac{\epsilon}{2}+\eta_{\frac{\epsilon}{2}}\cdot r_{\frac{\epsilon}{2}}=\epsilon$. Therefore, we give a polynomial-time upper bound for the mixing of double-loop Glauber on complete bipartite graph $G_{m,n}$:
Theorem S5.
@@
B Dense Bipartite Graphs
In this section, we extend the results from Supplementary Note A to non-complete bipartite graphs, providing proofs of Theorem 2 and Theorem 4. A natural approach is to leverage the aforementioned congestion calculations. We discuss dense bipartite graphs $G$: there exists a constant $\xi$, such that each vertex is non-adjacent to at most $\xi$ vertices in the opposite partition.
@@ B Dense Bipartite Graphs
[…]
we can rewrite (S30) as follows:
$\varrho:=$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{1}{\pi(M)P(M,M^{\prime})}\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\pi(I)\pi(F)|\gamma_{IF}^{(i)}|\bigg\}$ (S52)
$=$ $\max_{t=(M,M^{\prime})}\bigg\{|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\cdot\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}^{(i)}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\bigg\}.$ (S53)
For any bipartite graph $G$, let $G^{\prime}$ be the corresponding complete bipartite graph with the same vertex set as $G$. Let $\varrho_{t}^{\prime}$ denote the congestion of transition $t=(M,M^{\prime})$ in $G^{\prime}$, $E^{\prime}$ denote the edge set of $G^{\prime}$, $P^{\prime}(M,M^{\prime})$ denote the corresponding transition probability of double-loop Glauber for $G^{\prime}$, and $\operatorname{Haf}^{\prime}(M)$ denote the corresponding Hafnian of the subgraph induced by matching $M$ in $G^{\prime}$, which is equal to $|M|!$. We have already established an estimate for $\varrho_{t}^{\prime}$ in Supplementary Note A, and now we can use this estimate to bound $\varrho_{t}$ in $G$ by estimate $\operatorname{Haf}(M)$ via $\operatorname{Haf}^{\prime}(M)$ for any matching $M$.
@@
$1\geq\frac{\operatorname{Haf}(M)}{\operatorname{Haf}^{\prime}(M)}\geq\frac{(n^{\prime}-\xi)!}{n^{\prime}!}\geq\frac{1}{n^{\prime\xi}}\geq\frac{1}{n^{\xi}}.$ (S54)
@@
$\varrho=$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}(I)\lambda^{2|F|}\operatorname{Haf}(F)|\gamma_{IF}^{(i)}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}(\tilde{M})\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}(X)}\cdot|E|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg\}$ (S55)
$\leq$ $\max_{t=(M,M^{\prime})}\bigg\{\frac{\sum_{(I,F,i)\in cp(t)}p^{(i)}(I,F)\cdot\lambda^{2|I|}\operatorname{Haf}^{\prime}(I)\lambda^{2|F|}\operatorname{Haf}^{\prime}(F)|\gamma_{IF}^{(i)}|}{\lambda^{2|\tilde{M}|}\operatorname{Haf}^{\prime}(\tilde{M})n^{-\xi}\cdot\sum_{X\text{ is a matching}}\lambda^{2|X|}\operatorname{Haf}^{\prime}(X)n^{-\xi}}\cdot|E^{\prime}|\frac{1+\lambda^{2}}{{\lambda}^{2}}\bigg\}$ (S56)
$\leq$ $n^{2\xi}\varrho^{\prime}.$ (S57)
@@
By the analysis of Lemma S6, we have the congestion $\varrho^{\prime}$ of double-loop Glauber on complete bipartite graph $G_{m,n}$ is bounded by $\varrho^{\prime}=8mn^{4}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}$. Then we have the congestion $\varrho$ of double-loop Glauber on bipartite graph $G$ is bounded by $\varrho\leq n^{2\xi}\cdot\varrho^{\prime}=8mn^{4+2\xi}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}$. Thus again by Theorem S2, the mixing time of double-loop Glauber on complete bipartite graph $G_{m,n}$ is less than $16mn^{4+2\xi}\cdot\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\cdot\left(\ln{\epsilon^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)$. ∎
Furthermore, we complete the proof of Theorem 4 by extending the analysis of Theorem S5 to the case of dense bipartite graphs.
@@
$\eta_{\frac{\epsilon}{2}}=\frac{\epsilon}{2}\left(16mn^{4+2\xi}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)\right)^{-1}$ (S58)
in Lemma 1. Then the failure probability of each inner layer sampling step can be bounded within $\eta_{\frac{\epsilon}{2}}$. Again by Lemma S6, after
$16mn^{4+2\xi}\frac{(\lambda^{2}+1)(8\lambda^{2}-1)}{4\lambda^{2}-1}\left(\ln{{\epsilon}^{-1}}+2n\ln|{\lambda}|+2m\ln{m}\right)$ (S59)